The javascript on this page is used to preload the images for links and to swap the images used for links
go to content
layout image go to introduction
go to science datago to mission status reportgo to location / positiongo to 3d data mapsgo to moon VRMLgo to further reading
layout image
Layout imagego to home
go to projectgo to resultsGo to datavisgo to historyGo to scienceGo to educationGo to archivesgo to resources


Lunar Prospector Status Report #3

January 8, 1998 - 12:00 noon EST (9:00 a.m. PST)

Lunar Prospector mission operations personnel report the successful completion of activities on day 2 of the mission timeline. Excellent support from the Deep Space Network has ensured no significant outages in telemetry or command uplink. The monitoring of engineering and science data was the primary activity during this first full-day of spacecraft operation. All systems appear normal.

Between 1:09 a.m. and 2:44 a.m. (EST), the Electron Reflectometer, Neutron Spectrometer, and Gamma Ray Spectrometer instruments were commanded to high voltage. There were no anomalies and all instruments are currently providing good data. The current focus for the instruments is to collect calibration data enroute to the moon.

The second trajectory correction maneuver (TCM) was performed between 3:25 and 3:40 a.m. (EST). It was relatively small, only 8.4 m/s, and the navigation team was busy verifying the results during the overnight session. A third TCM has been scheduled for Friday night/Saturday morning to trim out any final errors.

It had been scheduled that the spacecraft would be placed into its attitude for lunar orbit insertion (LOI) burn after the TCM tonight, but that has been postponed until tomorrow afternoon. This will allow the dynamics engineer to get an update on the initial attitude, which changed slightly due to the TCM.

The current configuration of the spacecraft is as follows: it is spinning at 12.685 rpm, at cruise attitude, and all instruments are on. Data is currently being transmitted to the ground tracking station at 3600 bps (science format) on the omni antenna.

Activities for the next 24 hours include a command session beginning about 8 p.m. EST (Jan. 8) to command two of the instruments which may need gain changes, and to reorient the spacecraft to the LOI attitude.

David Morse
Ames Research Center
Moffett Field, CA 94035